# Transforming Policy-Car Swerving for Mitigating Stop-and-Go Traffic Waves: A Practice-Oriented Jam-Absorption Driving Strategy

### v1

@@ Abstract
Stop-and-go waves, as a major form of freeway traffic congestion, cause severe and long-lasting adverse effects, including reduced traffic efficiency, increased driving risks, and higher vehicle emissions. Amongst the highway traffic management strategies, jam-absorption driving (JAD), in which a dedicated vehicle performs “slow-in” and “fast-out” maneuvers before being captured by a stop-and-go wave, has been proposed as a potential method for preventing the propagation of such waves. However, most existing JAD strategies remain impractical mainly due to the lack of discussion regarding implementation vehicles and operational conditions. Inspired by real-world observations of police-car swerving behavior, this paper first introduces a Single-Vehicle Two-Detector Jam-Absorption Driving (SVDD-JAD) problem, and then proposes a practical JAD strategy that transforms such behavior into a maneuver capable of suppressing the propagation of an isolated stop-and-go wave. Five key parameters that significantly affect the proposed strategy, namely, JAD speed, inflow traffic speed, wave width, wave speed, and in-wave speed, are identified and systematically analyzed. Using a SUMO-based simulation as an illustrative example, we further demonstrate how these parameters can be measured in practice with two stationary roadside traffic detectors. The results show that the proposed JAD strategy successfully suppresses the propagation of a stop-and-go wave, without triggering a secondary wave. This paper is expected to take a significant step toward making JAD practical, advancing it from a theoretical concept to a feasible and implementable strategy. To promote reproducibility in the transportation domain, we have also open-sourced all the code on our GitHub repository (https://github.com/gotrafficgo).
@@ I Introduction
Stop-and-go waves (also referred to as wide moving jams or traffic oscillations) are a common form of traffic congestion on freeways worldwide, as empirically illustrated in Figure 1. Such a traffic pattern forces vehicles to repeatedly accelerate and decelerate, resulting in reduced traffic efficiency, longer travel times [33, 34], higher fuel consumption and emissions [21, 17], and an increased risk of traffic accidents [30, 32].
Figure 1: Global observations on freeway stop-and-go waves [9]. The striped low-speed regions propagating upstream against the direction of traffic flow are referred to as stop-and-go waves. (a) Nashville, TN, USA: 6.7 km (sourced from [6]); (b) Osaka City, Japan: 2 km (sourced from [5]); (c) California, USA: 0.6 km (sourced from [10]); (d) Rotterdam, Netherlands: 23 km (sourced from [7]); (e) Florida, USA: 190 km (sourced from [28]); (f) London, UK: 18 km (sourced from [26]).
@@
Amongst the highway traffic management strategies, jam-absorption driving (JAD), in which a dedicated vehicle performs “slow-in” and “fast-out” maneuvers before being captured by a stop-and-go wave, has been proposed as a potential method for preventing the propagation of such waves [2, 1]. Early studies focused on theoretical questions: Can JAD suppress stop-and-go waves? And will secondary waves be triggered? Initial analyses, based on kinematic wave theory (KWT) and fundamental diagram (FD) shapes, concluded that JAD is effective [24]. Later, microscopic models such as the Helly CF model and IDM were used to trace vehicle dynamics more explicitly [29, 25], though their conclusions rely on strong model assumptions. From 2017 onward, more practical, implementation-oriented JAD strategies have been proposed [12]. Core tasks include determining when to activate and deactivate a JAD vehicle and selecting the appropriate vehicle. The following works have incorporated wave propagation prediction [12], influential subspace concepts [15], capacity-drop considerations [8], multiple-wave suppression [22], and traffic efficiency and safety optimization [35, 31, 8]. Simulation is also employed to demonstrate how JAD vehicles can effectively increase flow and suppress waves.
@@
For a more detailed introduction and discussion of recent progress, one may refer to our latest review article [9]. Generally speaking, although JAD strategies have been extensively studied, most of these studies assume that connected and automated vehicles (CAVs) are available to perform the JAD tasks, and that complete traffic information is fully accessible. Even with the rapid development of CAV technology, fully controlling one or a few CAVs in real-world traffic to implement actual traffic management remains highly challenging. Moreover, the issue of JAD triggering secondary waves has been raised in nearly all existing studies, yet few have actually addressed it. Therefore, while JAD is theoretically appealing, its practical implementation is still far from realization.
@@
Occasionally, we observed a type of police-car swerving behavior frequently appearing on social media. As shown in Figure 2 and the footage in [3], a police car was recorded maneuvering at high speed across all lanes on a freeway in California, USA, to control the following traffic. This real-world behavior demonstrates the practical feasibility of deploying a slow-moving JAD vehicle to moderate the speed of inflow. Although the police car’s goal was still unclear to the authors (not jam absorption at least, we believe), it provides strong inspiration for potential real-world implementation of JAD.
@@
Figure 2: Police car swerving at high speed on a freeway, California, USA (Snapshots from the footage in [3]).
@@
Building on this real-world observation, this paper aims to develop a practically feasible strategy for suppressing stop-and-go waves. To this end, we formulate a Single-Vehicle Two-Detector Jam-Absorption Driving (SVDD-JAD) problem, in which only two stationary roadside traffic detectors (upstream and downstream) and a single JAD vehicle are assumed to be available. The objective is to block the propagation of a stop-and-go wave, specifically an isolated wave, distinct from the periodic stop-and-go waves associated with fixed-location bottlenecks, under these minimal conditions.
@@
We further conduct a detailed analysis of this problem, propose a corresponding JAD strategy, and identify five key parameters that significantly influence its performance: JAD speed, inflow traffic speed, wave width, wave speed, and in-wave speed. The effects of these parameters on the strategy are thoroughly analyzed. Using SUMO-based simulation, we discuss how these parameters can be measured in practice with only two stationary traffic detectors, and demonstrate the effectiveness of the proposed JAD strategy. The triggering of secondary waves is successfully avoided by incorporating traffic stability considerations into the selection of the JAD speed. By leveraging police-car swerving behavior to implement JAD, we believe this work substantially advances JAD from a purely theoretical concept toward practical application. Moreover, to promote reproducibility in the transportation domain, we have open-sourced all the code on our GitHub repository (https://github.com/gotrafficgo).
@@
The remainder of the paper is organized as follows. Section II introduces the SVDD-JAD problem. Section III proposes the JAD strategy and provides solutions to several key questions. Section IV presents a thorough theoretical analysis of the key parameters and demonstrates the results with numerical experiments. Section V uses a SUMO-based simulation to illustrate how these parameters can be measured in practice with only two stationary roadside traffic detectors and to validate the effectiveness of the proposed strategy. Finally, conclusions are drawn in Section VI.
@@ II Problem Statement
Inspired by the real-world observation of policy-car swerving on the freeway in Figure 2 and the footage in [3], we introduce an SVDD-JAD problem as follows.
@@
Single-Vehicle Two-Detector Jam-Absorption Driving (SVDD-JAD)
@@
Note that the targeted stop-and-go wave is an isolated wave, as illustrated by the wave labeled “A jam wave” in Figure 1. Such waves may originate from traffic disturbances under high-density conditions. Many existing variable speed limit (VSL) strategies, such as the well-known SPECIALIST approach [14, 13], are designed to address these types of waves. In contrast, the proposed strategy does not aim to resolve periodic stop-and-go waves associated with fixed-location bottlenecks. Many existing JAD studies seem not carefully clarify this.
Figure 3 describes the problem, in which $x_{A}$ denote the location of the stand-by dedicated vehicle for JAD; $x_{\mathrm{d}}$ and $x_{\mathrm{u}}$ are the locations of the downstream and upstream detectors, respectively; $w$ denote the stop-and-go wave speed; $v_{\mathrm{t}}$ and $v^{*}$ are the car-following speed and the pre-determined JAD speed of the JAD vehicle, respectively.
@@
  • $E(t_{E},x_{E})$ is the point where the back of a stop-and-go wave is detected by the downstream detector at location $x_{\mathrm{d}}$.
  • $F(t_{F},x_{F})$ is the point where the front of a stop-and-go wave is detected by the downstream detector at location $x_{\mathrm{d}}$.
@@
where $A\xrightarrow{v_{\mathrm{t}}}B$ is a process of floating within traffic and waiting for an appropriate moment, and $B\xrightarrow{v^{*}}C$ is the actual JAD process.
@@
$v_{\mathrm{t}}>v^{*}>v_{\mathrm{w}}.$ (2)
Points $E$ and $F$ are associated with the stop-and-go wave detected by the downstream detector at $x_{\mathrm{d}}$. The temporal width of the stop-and-go wave can be denoted by $\Delta_{\mathrm{w}}=t_{E}-t_{F}$. In addition, it is clear to see from Figure 3 that $x_{\mathrm{d}}=x_{E}=x_{F}$.
The key questions to solve the SVDD-JAD problem are as follows.
@@
It is worth noting that terminating the JAD task earlier or later than point $C$ (e.g., at $C^{\prime}$ or $C^{\prime\prime}$) will lead to collision or inefficiency, as shown in Figure 3.
@@
    Stop-and-go waves propagate backward at an approximately constant speed. Actually, this assumption is consistent with real-world observations[23, 19, 36, 18, 11].
@@
    The length of the backward-moving queue is also approximately constant, instead of continuously increasing, as shown in the real-world observations such as those in Figure 1. If the queue can dissipate itself (e.g., due to the inflow decrease), there is no need to dispatch a JAD vehicle. This assumption can be relaxed by installing one more downstream detector and estimating the rate of wave growth.
@@
    In the theoretical analysis, we assume an instantaneous speed change in the analysis, which simplifies the analysis. In simulation or practice, a gradually-changing speed would not largely change the results, as shown in [12].
@@ III-A Q1: What is the JAD speed?
First, the JAD speed should be lower than the ahead-traffic speed, i.e., $v^{*}<v_{\mathrm{t}}$, according to the original “slow-in” and “fast-out” idea of JAD.
@@
According to fundamental traffic flow theory, secondary waves or traffic breakdowns are primarily associated with traffic instability. Generally speaking, the low density is stable while the high density is easy to break down [^1] The notion of stability can differ depending on the perspective. From a microscopic or dynamical viewpoint, a high-density flow is linearly unstable, i.e., small perturbations can grow into stop-and-go waves. However, from a macroscopic or phase perspective, once traffic has already entered a congested state, the flow pattern becomes relatively stable and does not further break down. In this study, the instability we refer to is of the first type, i.e., from a microscopic viewpoint. (Figure 4). There might be a critical density ($k^{\prime}$) and the associated critical speed ($v^{\prime}$), which separates the states [^2] The critical density and speed here may differ from those commonly used in traffic flow theory to separate free-flow and congested conditions.. Therefore, to avoid triggering the secondary wave, the JAD speed should be higher than the critical speed, i.e., $v^{*}\geq v^{\prime}$.
@@ III-B Q2: What is the JAD plan? Part I: Points $B$ and $C$
  • The leading vehicle (or traffic) speed $v_{\mathrm{t}}$,
  • The back and front of a stop-and-go wave detected by the downstream detector at $x_{\mathrm{d}}$, i.e., points $E$ and $F$,
@@ III-C Q2: What is the JAD plan? Part II: Point AA
In practice, the location and moment (i.e., point $A$) to dispatch a JAD vehicle may be determined by traffic managers: A JAD vehicle initially stands by at a fixed roadside location or an on-ramp, denoted by $x_{A}$. Once a downstream stop-and-go wave is confirmed to be developing, the JAD vehicle is dispatched at time $t_{A}$, subject to practical considerations such as the emergence of a sufficiently large traffic gap. However, this appointment is also constrained by a feasible range of points $A$, and this subsection focuses on determining the feasible range of the starting point $A$.
With the aid of Figure 5, we analyze the feasible region of point $A$. We begin by ignoring point AA [^3] In Section III-C, point $A$ is assumed to be known, and the corresponding JAD plan is then constructed. This sequence is consistent with the natural decision-making process in practice, where a JAD opportunity is first identified and subsequently planned. Here, however, in order to characterize the feasible region of point $A$, we do not assume that point $A$ is known a priori. and then consider the JAD Triangle $BCD$, whose sides $BC$, BDBD, and $CD$ have fixed slopes $^{*}, vv^{*}$, and $v_{\mathrm{w}}$, respectively. Points $C$ and $D$ are constrained to lie on the parallel lines $CE$ and $DF$, both with slope $w$. Since the shape of the JAD Triangle is fixed and segment $CD$ is confined between these two parallel lines, the triangle $BCD$ can only undergo rigid translations along a direction with slope $w$, under which point $B$ traces a straight line, denoted by $\ell$, with slope $w$.
Figure 5: Feasible region of the dispatch time and location of a JAD vehicle (point AA).
The following conditions are given. Points $C$ lies on lines $CE$ in (4) and line $BC$ in (9), giving
@@
From a reactive perspective (i.e., without prediction), a JAD vehicle can be dispatched only after the formation of a stop-and-go wave is confirmed at time $t_{E}$ ($t_{E}>t_{F}$); see Case 1 in Figure 5. The upstream detector located at $x_{\mathrm{u}}$ defines the earliest admissible stand-by location of the JAD vehicle; see Case 2 in Figure 5. Therefore, the temporal and spatial lower bounds are given as follows:
@@
These three points form a triangle, which defines the feasible region of point $A$, since point $A$ is the backward extension of point $B$ along the space–time direction.

### v2

@@ Abstract
Stop-and-go traffic waves, as a major form of freeway traffic congestion, cause severe and long-lasting adverse effects, including reduced traffic efficiency, increased driving risks, and higher vehicle emissions. Amongst the freeway traffic management strategies, jam-absorption driving (JAD), in which a dedicated vehicle performs “slow-in” and “fast-out” maneuvers before being captured by a stop-and-go wave, has been proposed as a potential method for preventing the propagation of such waves. However, most existing JAD strategies remain impractical mainly due to the lack of discussion regarding implementation vehicles and operational conditions. Inspired by real-world observations of police-car swerving behavior, this paper first introduces a Single-Vehicle Double-Detector Jam-Absorption Driving (SD-JAD) problem, and then proposes a practical JAD strategy that transforms such behavior into a JAD strategy capable of suppressing the propagation of an isolated stop-and-go wave. Five key parameters that significantly affect the proposed strategy, namely, JAD speed, inflow traffic speed, wave width, wave speed, and in-wave speed, are identified and systematically analyzed. Using a SUMO-based simulation as an illustrative example, we further demonstrate how these parameters can be measured in practice with two stationary roadside traffic detectors available. The results show that the proposed JAD strategy successfully suppresses the propagation of a stop-and-go wave, particularly with no secondary waves triggered. This paper is expected to take a significant step toward making JAD practical, advancing it from a theoretical concept to a feasible and implementable strategy. To promote reproducibility in the transportation domain, we have also open-sourced all the code on our GitHub repository (https://github.com/gotrafficgo).
@@ I Introduction
Stop-and-go traffic waves (also referred to as wide moving jams or traffic oscillations) are a common form of traffic congestion on freeways. Such a freeway traffic pattern forces vehicles to repeatedly accelerate and decelerate, resulting in reduced traffic efficiency, longer travel times [30, 31], higher fuel consumption and emissions [18, 14], and an increased risk of traffic accidents [27, 29]. Even worse, such stop-and-go structures can propagate and sustain themselves over tens of kilometers under sufficiently high traffic demand, which further exacerbates their negative impacts.
@@
Amongst the freeway traffic management strategies, jam-absorption driving (JAD), in which a dedicated vehicle performs “slow-in” and “fast-out” maneuvers before being captured by a stop-and-go wave, has been proposed as a potential method for preventing the propagation of such waves [2, 1]. Early studies focused on theoretical questions: Can JAD suppress stop-and-go waves? And will secondary waves be triggered? Initial analyses, based on the Kinematic Wave Theory and the Fundamental Diagram, concluded that JAD is effective [21]. Microscopic models such as the Helly car-following model and the Intelligent Driver Model were used to trace vehicle dynamics more explicitly [24, 22], though their conclusions rely on strong model assumptions. From 2017 onward, more practical, implementation-oriented JAD strategies have been proposed [8]. Core tasks include determining when to activate and deactivate a JAD vehicle and selecting the appropriate vehicle. The following works have incorporated wave propagation prediction [8], influential subspace concepts [12], capacity-drop considerations [5], multiple-wave suppression [19], and traffic efficiency and safety optimization [32, 28, 5]. Simulation is typically employed to demonstrate how JAD vehicles can effectively increase flow and suppress waves. For a more detailed introduction and discussion of recent JAD progress, one may refer to our latest review article [6].
@@
Generally speaking, although JAD strategies have been extensively studied, most assume the availability of Connected and Automated Vehicles (CAVs) to perform JAD tasks and full access to traffic information to determine a specific JAD plan. Unfortunately, even with the rapid development of CAV technology, fully controlling one or a few CAVs in real-world traffic to implement actual traffic management remains highly challenging and rarely seen. Moreover, the issue of JAD triggering secondary waves has been raised in nearly all existing studies, yet few have actually addressed it. Therefore, while JAD is theoretically appealing, its practical implementation is still far from realization.
@@
Occasionally, we observed a type of police-car swerving behavior frequently appearing on social media. As shown in Figure 1 and the footage in [3], a police car was recorded maneuvering at high speed across all lanes on a freeway in California, USA, to control the following traffic. Although the police car’s goal was still unclear to the authors (not jam absorption at least, we believe), this real-world behavior demonstrates the practical feasibility of deploying a slow-moving vehicle to moderate the speed of inflow.
@@
Figure 1: Police-car swerving at high speed on a freeway, California, USA (Snapshots from the footage in [3]).
@@
Building on this real-world observation, this paper aims to develop a practically feasible strategy for suppressing stop-and-go waves. To this end, we formulate a Single-Vehicle Double-Detector Jam-Absorption Driving (SD-JAD) problem, in which only two stationary roadside traffic detectors (upstream and downstream) and a single JAD vehicle are assumed to be available. Under these minimal conditions, the JAD objective is to block the propagation of a stop-and-go wave, specifically an isolated wave, which is distinct from the periodic stop-and-go waves associated with fixed-location bottlenecks.
@@
We further propose a corresponding JAD strategy and identify five key parameters that significantly influence its performance, namely, JAD speed, inflow traffic speed, wave width, wave speed, and in-wave speed. The effects of these parameters on the strategy are thoroughly analyzed. Using SUMO-based simulation, we discuss how these parameters can be measured in practice with only two stationary traffic detectors, and demonstrate the effectiveness of the proposed JAD strategy. The triggering of secondary waves is successfully prevented by taking traffic stability into account when selecting the JAD speed. By leveraging police-car swerving behavior to implement JAD, we believe this work substantially advances JAD from a purely theoretical concept toward practical application. Moreover, to promote reproducibility in the transportation domain, we have open-sourced all the code on our GitHub repository (https://github.com/gotrafficgo).
@@
The remainder of the paper is organized as follows. Section II introduces the SD-JAD problem. Section III proposes the JAD strategy and provides solutions to several key questions. Section IV presents a thorough theoretical analysis of the key parameters and demonstrates the results with numerical experiments. Section V uses a SUMO-based simulation to illustrate how these parameters can be measured in practice with only two stationary roadside traffic detectors and to validate the effectiveness of the proposed strategy. Finally, conclusions are drawn in Section VI.
@@ II Problem Statement
Inspired by the real-world observation of policy-car swerving on the freeway in Figure 1 and the footage in [3], we introduce an SD-JAD problem as follows.
@@
Single-Vehicle Double-Detector Jam-Absorption Driving (SD-JAD)
@@
Note that the targeted stop-and-go wave is an isolated wave with high-density free-flow conditions both upstream and downstream. Such waves may originate from disturbances under high traffic density. Many existing variable speed limit (VSL) strategies, such as the well-known SPECIALIST [10, 9], are designed to address these types of waves [6]. In contrast, the proposed strategy does not aim to resolve periodic stop-and-go waves associated with a fixed-location bottleneck. Many existing JAD studies seem not carefully clarify this.
Figure 2 describes the problem, in which $x_{A}$ denotes the location of the stand-by dedicated vehicle for JAD; $x_{\mathrm{d}}$ and $x_{\mathrm{u}}$ are the locations of the downstream and upstream detectors, respectively; $w$ denotes the stop-and-go wave speed; $v_{\mathrm{t}}$ and $v^{*}$ are the car-following speed and the pre-determined JAD speed of the JAD vehicle, respectively.
@@
  • $E(t_{E},x_{E})$ is the point where the temporal back of a stop-and-go wave is detected by the downstream detector at location $x_{\mathrm{d}}$.
  • $F(t_{F},x_{F})$ is the point where the temporal front of a stop-and-go wave is detected by the downstream detector at location $x_{\mathrm{d}}$.
@@
where $A\xrightarrow{v_{\mathrm{t}}}B$ is a process of floating within traffic and waiting for an appropriate moment and $B\xrightarrow{v^{*}}C$ is the actual JAD process.
@@
$v_{\mathrm{t}}>v^{*}>v_{\mathrm{w}}\geq 0.$ (2)
Points $E$ and $F$ are associated with the stop-and-go wave detected by the downstream detector at $x_{\mathrm{d}}$, and thus $x_{\mathrm{d}}=x_{E}=x_{F}$. The temporal width of the stop-and-go wave can be denoted by $\Delta_{\mathrm{w}}=t_{E}-t_{F}$.
The key questions to solve the SD-JAD problem are as follows.
@@
It is worth noting that terminating the JAD task earlier or later than point $C$ (e.g., at $C^{\prime}$ or $C^{\prime\prime}$) will lead to danger or inefficiency, as shown in Figure 2.
@@
    Stop-and-go waves propagate backward at an approximately constant speed, which is quite consistent with real-world observations[20, 16, 33, 15, 7].
@@
    The length of the backward-moving queue is also approximately constant, instead of continuously increasing, as shown in the real-world observations such as those in Figure LABEL:fig:wave. If the queue can shrink itself (e.g., due to the inflow decrease), there is no need to dispatch a JAD vehicle. This assumption can be relaxed by incorporating other information to estimate the changes of wave width.
@@
    For simplicity, we assume an instantaneous speed change in the theoretical analysis. In simulation or practice, a gradually-changing speed would not largely change the results, as shown in [8].
@@ III-A Q1: What is the JAD speed?
First, the JAD speed should be lower than the speed of the surrounding or upcoming traffic, i.e., $v^{*}<v_{\mathrm{t}}$, according to the original “slow-in” and “fast-out” idea of JAD.
@@
According to fundamental traffic flow theory, secondary waves or traffic breakdowns are primarily associated with traffic instability. Generally speaking, the low density is stable while the high density is easy to break down [^1] The notion of stability can differ depending on the perspective. From a microscopic or dynamical viewpoint, a high-density flow is linearly unstable, i.e., small perturbations can grow into stop-and-go waves. However, from a macroscopic or phase perspective, once traffic has already entered a congested state, the flow pattern becomes relatively stable and does not further break down. In this study, the instability we refer to is of the first type, i.e., from a microscopic viewpoint. (Figure 3). There might be a critical density (denoted by $k^{\prime}$) and the associated critical speed (denoted by $v^{\prime}$), which separates the states [^2] The critical density and speed here may differ from those commonly used in traffic flow theory to separate free-flow and congested conditions.. Therefore, to avoid triggering the secondary wave, the JAD speed should be higher than the critical speed, i.e., $v^{*}\geq v^{\prime}$.
@@ III-B Q2: What is the JAD plan? Part I: Points $B$ and $C$
  • The surrounding or upcoming traffic speed $v_{\mathrm{t}}$,
  • The temporal back and front of a stop-and-go wave at a fixed location, detected by the downstream detector at $x_{\mathrm{d}}$, i.e., points $E$ and $F$,
@@ III-C Q2: What is the JAD plan? Part II: Point AA
In practice, the location and moment (i.e., point $A$) to dispatch a JAD vehicle may be determined by traffic managers. Specifically, a JAD vehicle initially stands by at a fixed roadside location or an on-ramp, denoted by $x_{A}$. Once a downstream stop-and-go wave is confirmed to be developing, the JAD vehicle is dispatched at time $t_{A}$, subject to practical considerations such as the emergence of a sufficiently large traffic gap. However, this appointment is also constrained by a feasible range of points $A$, and this subsection focuses on determining the feasible range of the starting point $A$.
With the aid of Figure 4, we analyze the feasible region of point $A$. We begin by ignoring point AA [^3] In Section III-C, point $A$ is assumed to be known, and the corresponding JAD plan is then constructed. This sequence is consistent with the natural decision-making process in practice, where a JAD opportunity is first identified and subsequently planned. Here, however, in order to characterize the feasible region of point $A$, we do not assume that point $A$ is known a priori. and considering the JAD Triangle $BCD$, whose sides $BC$, BDBD, and $CD$ have fixed slopes $v^{*}$, $v_{\mathrm{t}}$, and $v_{\mathrm{w}}$, respectively. Points $C$ and $D$ are constrained to lie on the parallel lines $CE$ and $DF$, both with slope $w$. Since the shape of the JAD Triangle $BCD$ is fixed and segment $CD$ is confined between these two parallel lines, the JAD Triangle $BCD$ can only undergo rigid translations along a direction with slope $w$, under which point $B$ traces a straight line, denoted by $\ell$, with slope $w$.
Figure 4: Feasible region of the dispatch time and location (point AA) of a JAD vehicle.
We have the following conditions. Points $C$ lies on lines $CE$ in (4) and line $BC$ in (9), giving
@@
From a reactive perspective (i.e., without prediction), a JAD vehicle can be dispatched only after the formation of a stop-and-go wave is confirmed at time $t_{E}$ ($t_{E}>t_{F}$); see Case 1 in Figure 4. The upstream detector located at $x_{\mathrm{u}}$ defines the farthest admissible stand-by location of the JAD vehicle; see Case 2 in Figure 4. Therefore, the temporal and spatial lower bounds are given as follows:
@@
These three points form a triangle, which defines the feasible region of point $A$, since point $A$ is the backward extension of point $B$ along the space-time direction.